OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form

OpenStudy (anonymous):

\[-x+3 y-8 = 0\]

Directrix (directrix):

1x - 3y = -8

OpenStudy (anonymous):

-x+3y-8=0 is what i got

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. http://jwilson.coe.uga.edu/emt668/emat6680.2002/jackson/chapter%205%20lesson%20plan/day6.html

OpenStudy (anonymous):

m = 1/3 4 = (1/3)4 + b ->b = 4 - 4/3 = 8/3 Thus y = x/3 + 8/3 3y = x + 8 x - 3y = 8

OpenStudy (anonymous):

can some one do this one Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form

Directrix (directrix):

So, I should write: 1x + (-3) y = -8

OpenStudy (anonymous):

y -5 = 3 ( x -2) y = 3x -1

OpenStudy (anonymous):

yahooyc, did you see my reply?

OpenStudy (anonymous):

i did but i dont have that option in my awser key this is what i got y – 2 = 3(x – 5) y + 8 = 3(x + 5) y + 5 = 3(x + 2) y – 5 = 3(x – 2)

OpenStudy (anonymous):

i dont know which one to pick

OpenStudy (anonymous):

y – 5 = 3(x – 2)

OpenStudy (anonymous):

The first step of my answer!

OpenStudy (anonymous):

thank you everyone

OpenStudy (anonymous):

The equation through 2 pts with slope is: y - yA = m ( x - xA)

OpenStudy (anonymous):

I need to fix my answer! 3y = x + 8 x - 3y = -8